OpenStudy (deepolisnoob):

Is sin^(-1)cosx equal to 1?

OpenStudy (misty1212):

HI!!

OpenStudy (deepolisnoob):

\[\sin^{-1}\cos(x)\]

OpenStudy (deepolisnoob):

Hey there!

OpenStudy (misty1212):

\[\sin^{-1}(\cos(x))\] like that?

OpenStudy (deepolisnoob):

Yeah, that is it. I'm thinking it is one since cos and sin are integrals/derivatives of each other, and they both equal one in the double angle identity

OpenStudy (misty1212):

no , pretty sure it is not one

OpenStudy (misty1212):

i believe it is some kind of saw tooth function

OpenStudy (misty1212):

here is a nice picture http://www.wolframalpha.com/input/?i=arcsin(cos(x))

OpenStudy (deepolisnoob):

mmm, okay. The part where this comes from is this other equation... \[\int\limits_{cosx}^{0}\frac{ dt }{ \sqrt{1-t^2} }\]

OpenStudy (misty1212):

hard to say what it is really \[\sin^{-1}(\cos(0))=\sin^{-1}(1)=\frac{\pi}{2}\]for example

OpenStudy (deepolisnoob):

mmm, yeah, I figured. I was trying to see if maybe they would cancel each other in some way through algebra

OpenStudy (misty1212):

can you post the entire quesition?

OpenStudy (misty1212):

at this point you get exactly what you wrote as an answer, a function, not a number

OpenStudy (misty1212):

\[\sin^{-1}(0)-\sin^{-1}(\cos(x))\]

OpenStudy (deepolisnoob):

The question is: Find dy/dx of the following equation: \[y=\int\limits_{cosx}^{0}\frac{ dt }{ \sqrt{1-t^2} }\]

OpenStudy (misty1212):

aka \[-\sin^{-1}(\cos(x))\]

OpenStudy (deepolisnoob):

they dont give me much to work with

OpenStudy (misty1212):

lol i knew it !!

OpenStudy (misty1212):

you are doing too much work

OpenStudy (misty1212):

the fundamental theorem of calculus says the derivative of the integral is the integrand

OpenStudy (misty1212):

so first is \[\int_0^{\cos(x)}\frac{1}{\sqrt{1-t^2}}dt\]

OpenStudy (misty1212):

then replace \*t\) by \(\cos(x)\) and by the chain rule, multiply by the derivative of cosine

OpenStudy (misty1212):

oops meant first is \[-\int_0^{\cos(x)}\frac{1}{\sqrt{1-t^2}}dt\]

OpenStudy (misty1212):

then the derivative is \[-\frac{1}{\sqrt{1-\cos^2(x)}}\times (-\sin(x))\]

OpenStudy (misty1212):

you can clean it up a lot, like get rid of the two minus signs, and replace the denominator by \(|\sin(x)|\)

OpenStudy (deepolisnoob):

mmm, okay, thank you very much! I just didn't read the question and went straight into integrating lol. I really appreciate your help!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]